OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. 1. What property is shown by the equation? Column A Column B 1. (34 + 11) + 9 = (11 + 34) + 9 2. 25 + (–15 + 4) = (25 + (–15)) + 4 3. –(–21 + 8) = 21 + (–8) 4. –30 + (–16) = –16 + (–30) A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. 1. What property is shown by the equation? Column A Column B 1. (34 + 11) + 9 = (11 + 34) + 9 2. 25 + (–15 + 4) = (25 + (–15)) + 4 3. –(–21 + 8) = 21 + (–8) 4. –30 + (–16) = –16 + (–30)

OpenStudy (anonymous):

A. commutative property B. associative property C. opposite of a sum property